# Toward Metaphor-Fluid Conversation Design for Voice User Interfaces

### v1

@@ Abstract.
Metaphors play a critical role in shaping user experiences with Voice User Interfaces (VUIs), yet existing designs often rely on static, human-centric metaphors that fail to adapt to diverse contexts and user needs. This paper introduces Metaphor-Fluid Design, a novel approach that dynamically adjusts metaphorical representations based on conversational use-contexts. We compare this approach to a Default VUI, which characterizes the present implementation of commercial VUIs commonly designed around the persona of an assistant, offering a uniform interaction style across contexts. In Study 1 (N=130), metaphors were mapped to four key use-contexts—commands, information seeking, sociality, and error recovery—along the dimensions of formality and hierarchy, revealing distinct preferences for task-specific metaphorical designs. Study 2 (N=91) evaluates a Metaphor-Fluid VUI against a Default VUI, showing that the Metaphor-Fluid VUI enhances perceived intention to adopt, enjoyment, and likability by aligning better with user expectations for different contexts. However, individual differences in metaphor preferences highlight the need for personalization. These findings challenge the one-size-fits-all paradigm of VUI design and demonstrate the potential of Metaphor-Fluid Design to create more adaptive and engaging human-AI interactions.
@@ 1. Introduction
Voice User Interfaces (VUIs) are becoming an integral part of daily interactions, embedded in smart speakers, mobile devices, automobiles, and wearable technologies. As these systems proliferate, the ways in which users engage with and conceptualize them evolve dynamically. However, contemporary VUI design remains largely constrained by a single dominant metaphor—humanness (Doyle et al., 2019). Most commercial VUIs are designed to mimic human conversational abilities, using speech, turn-taking, and natural language processing to create an interaction model that feels intuitive (Edlund, 2019; Pradhan and Lazar, 2021). Within this overarching metaphor, VUIs are often framed as social entities that take on various roles, including teachers, therapists, coaches, advisors, and companions (Desai and Chin, 2023; Jung et al., 2019; Desai et al., 2023a; Wang et al., 2020; Desai et al., 2023b; Lee et al., 2021; Motalebi et al., 2019; Trajkova and Martin-Hammond, 2020; Purington et al., 2017). Among these, the assistant persona has emerged as the most prevalent, dominating commercial implementations and positioning the VUI as a helpful, subservient entity designed to execute user commands (Desai and Twidale, 2023; Doyle et al., 2021). While this metaphor provides a familiar interaction paradigm, it also imposes significant limitations. It assumes a fixed, one-size-fits-all identity for the VUI, failing to account for the fluid nature of human-VUI interactions across different tasks and contexts (Pradhan and Lazar, 2021; Desai et al., 2024a; Desai and Chin, 2023; Desai and Twidale, 2022). This rigidity leads to usability challenges (Cowan et al., 2017), mismatches between user expectations and system behavior (Doyle et al., 2021), and broader concerns surrounding anthropomorphic design choices (McMillan and Jaber, 2021; Desai and Twidale, 2022; Simpson and Crone, 2022; Pradhan and Lazar, 2021; Luger and Sellen, 2016; Turk, 2016).
@@
These challenges reflect a broader issue in interface design: the role of metaphors in shaping user interaction. Metaphors have played a foundational role in shaping Human-Computer Interaction (HCI) (Neale and Carroll, 1997; Laurel, 1997; Laurel and Mountford, 1990; Carroll and Thomas, 1982; Carroll et al., 1988; Don et al., 1992). In Graphical User Interfaces (GUIs), they are structured into primary and secondary layers: the desktop metaphor, for instance, serves as the primary organizing concept, with secondary metaphors—such as files, folders, and recycling bins—providing finer interaction affordances (Blackwell, 2006). These layered metaphors help users form coherent mental models of system functionality. In contrast, VUIs lack a stable visual representation, making their metaphorical framing less explicit. Instead, this framing is operationalized through persona design, where a VUI’s voice, linguistic style, and response behaviors collectively establish its identity (Pradhan and Lazar, 2021). Unlike GUIs, where metaphors map interactions onto familiar objects, VUIs inherently imply personhood, reinforcing anthropomorphic interpretations. The presence of human-like speech cues, names, and conversational norms encourages users to assign social roles to VUIs, often expecting them to behave with intelligence, adaptability, and even emotional awareness (Luger and Sellen, 2016; Sutton et al., 2019).
@@
Despite this anthropomorphic framing, users do not engage with VUIs as static entities (Pradhan et al., 2019). Instead, they shift their metaphorical framings depending on use-contexts—the functional settings in which interactions take place (Desai and Twidale, 2023). Prior research has identified several recurring use-contexts in VUI interactions, including commands, where users expect efficiency and compliance (e.g., ”Turn off the lights”); information seeking, where they seek structured knowledge retrieval (e.g., ”What is the capital of Japan?”); sociality, where they engage in casual conversation (e.g., ”Tell me a joke”); and error recovery, where they attempt to correct or refine system responses (e.g., ”That’s not what I meant—try again”) (Sciuto et al., 2018; Desai and Chin, 2023; Ammari et al., 2019; Bentley et al., 2018). In each of these use-contexts, users adopt distinct metaphorical descriptions: they conceptualize a VUI as a butler or secretary when issuing commands, as an expert librarian when retrieving information, and as a companion in social exchanges. However, commercial VUIs impose a singular, unchanging persona across all contexts, failing to adapt to the shifting expectations users bring to different types of interactions. This rigidity results in fundamental misalignments—when users anticipate authoritative expertise but receive generic, non-committal responses, or when they expect a social exchange but encounter the same neutral, task-oriented demeanor, the interaction feels unnatural and disengaging (Desai et al., 2023a). These breakdowns not only present usability challenges but also raise deeper ethical concerns.
@@
Building on prior work, we introduce Metaphor-Fluid Design, a novel design approach that moves beyond static persona models by enabling VUIs to dynamically shift their metaphorical framing based on conversational context. Rather than adhering to a singular metaphor—such as the assistant—this approach allows the system to adopt multiple personas, aligning its interaction style with the expectations of the task at hand. Crucially, Metaphor-Fluid Design also incorporates non-human and fictional metaphors, which have been shown to be both relevant and frequently used by users (Desai et al., 2024a; Desai and Twidale, 2023; Jung et al., 2022). While commercial VUIs default to anthropomorphic personas, users often describe them in non-human terms, likening them to machines, calculators, encyclopedias, or even mythical entities like genies or ghosts (Desai and Twidale, 2023). By broadening the metaphorical landscape beyond human roles, this approach creates greater flexibility in interaction design, allowing VUIs to respond more intuitively to diverse user expectations. This work is guided by two primary research questions:
@@
These questions are explored through two studies. The first study (N=130) maps user-generated metaphors to different conversational contexts, revealing systematic variations in metaphorical preferences across tasks. Findings show that users do not conceptualize VUIs through a singular metaphor but instead shift their framings based on the use-context. Specifically, users associate command-based interactions and error recovery with Guides (e.g., Genie, Search Engine), information-seeking tasks with Aides (e.g., Encyclopedia, ”Computer” from Star Trek), and social exchanges with Companions (e.g., Friend, Flmatmate). These patterns demonstrate that users instinctively align VUI metaphors with the demands of the task at hand, challenging the assumption that a single persona is sufficient for all interactions. The second study (N=91) empirically evaluates a Metaphor-Fluid VUI—which adapts its metaphorical framing dynamically—against a Default VUI that consistently operationalizes the assistant metaphor. Results indicate that Metaphor-Fluid design leads to significantly higher perceived likability, enjoyment, and intention to adopt, showing that aligning a VUI’s persona with context-specific expectations positively impacts user perceptions. However, findings also highlight individual differences in metaphor preferences, suggesting that while contextual adaptation enhances certain perceptions, incorporating personalization could further refine metaphor alignment for diverse users.
@@ 2.2. Metaphors in HCI
The application of metaphors in interface design extends beyond mere aesthetics, serving dual roles of familiarization and transportation, as outlined by Paul Heckel in his work on friendly software design (Heckel, 1984). Familiarization employs metaphors to introduce recognizable concepts, facilitating user comprehension and mental model construction. Transportation, on the other hand, utilizes these mental models to create immersive experiences. In the context of VUIs, the concept of humanness frequently serves as the primary metaphor, acting as a conduit for familiarization and transportation effects.
@@ 2.3. Metaphors in VUIs
Epley et al.’s influential three-factor theory of anthropomorphism explains why and when humans attribute human-like characteristics to non-human agents. The factors—elicited agent knowledge (drawing on knowledge of human behavior), effectance motivation (the need to predict and understand the agent’s actions), and sociality motivation (seeking social connection)—are key to understanding why users often interact socially with VUIs (Epley et al., 2007). The ”humanness” metaphor leverages users’ natural inclination to perceive technology through a social lens, making VUIs feel more engaging and personable. As a result, these systems incorporate anthropomorphic cues like natural language processing and human-like voice modulation to reinforce the perception of social presence and competence (Pradhan and Lazar, 2021; Purington et al., 2017).
@@
    Simplistic Representations: While users develop complex and context-dependent mental models of VUIs that change with the type of interaction or domain of use, designers often rely on a single overarching metaphor to represent these systems. This singular metaphorical framing may fail to capture the nuanced ways in which users interact with VUIs across different contexts (Desai and Twidale, 2022). For example, in a user study involving older adults interacting with a VUI designed as an ”exercise coach,” participants expressed a preference for the VUI to also behave like a ”friend” when initiating the session, checking on their progress, and providing motivation (Desai et al., 2023a). The complexity of user expectations and the diverse roles VUIs play are not adequately addressed by a one-size-fits-all metaphor, which can lead to a mismatch between what users expect and what the system delivers, ultimately impacting usability and user satisfaction. By providing only a simplistic metaphor, designers risk oversimplifying the rich, multi-faceted experiences users have with VUIs, thereby limiting the potential for these systems to meet diverse user needs effectively. Moreover, users’ perceptions of VUIs are also highly fluid, shifting depending on the context and specific interactions. Pradhan et al. (Pradhan and Lazar, 2021) found that users alternately view VUIs as either a ”person” or a ”thing,” or something in between, depending on the specific task or setting.
@@
    Misaligned Mental Models: Misaligned mental models arise when the human-like metaphors underlying VUIs heighten user expectations, leading to either over or under-calibration of trust. Users may overestimate the VUI’s abilities, assuming it possesses human-like cognitive and conversational skills, or underestimate its utility due to skepticism about its capabilities. (Cowan et al., 2017) This misalignment of expectations impacts trust calibration—users may become overly reliant on the VUI, expecting it to perform beyond its technical limitations, or conversely, they may dismiss its utility prematurely (Luger and Sellen, 2016). These fluctuating expectations can result in inconsistent or infrequent use (Cowan et al., 2017) and, eventually, complete abandonment of the technology (Trajkova and Martin-Hammond, 2020). The inconsistency in how users perceive the capabilities of VUIs, especially when the system fails to meet expectations in specific scenarios, resulting in conversational errors, further exacerbates the misalignment and undermines user trust (Baughan et al., 2023).
The evolving landscape of VUI design reflects a broader understanding of the complexities involved in metaphor selection. While a single human metaphor offers an accessible and engaging way for users to interact with technology, its limitations require thoughtful application to prevent unintended ethical and usability issues.
2.4. Mitigating Challenges in VUI Design
@@
As the VUI design space evolves, various approaches have been proposed to address the challenges surrounding VUI persona design. One promising approach involves developing VUIs in non-human roles. Desai and Twidale (Desai and Twidale, 2023), through a literature review and user study, identified how metaphors are used by users, designers, manufacturers, and researchers in relation to VUIs. Based on their analysis, they developed a framework for metaphor contextualization in VUIs, categorizing metaphor use along four dimensions: type (human, non-human, fictional), mentioned by (users, designers, manufacturers, researchers), the knowledge it draws upon, and the motivation behind its use. Their findings revealed that while human metaphors dominate VUI design and research, non-human and fictional metaphors were also popular, especially among users for sense-making. This contrast highlights a dichotomy between how designers and researchers present VUIs and how users interpret them. This observation points to the need for a ”pragmatic approach” (Carroll et al., 1988) to systematically explore the design space between system capabilities and user expectations.
@@
In the broader CUI literature, there have been recent notable advancements in exploring the feasibility of designing conversational agents with non-human roles. Jung et al. (Jung et al., 2022), for instance, used the Great Chain of Being (GCOB) framework [^2] A theoretical framework that organizes the world into a hierarchical structure using metaphors of God, humans, animals, plants, and inanimate objects. to investigate how non-human metaphors (such as God, Animal, Plant, and Book) compared to human metaphors influence user perceptions of chatbots. Their results showed that the human metaphor did not significantly alter chatbot evaluations compared to non-human metaphors. Similarly, in a related study involving VUIs, researchers examined human versus non-human metaphors in the health and finance domains. They found that in the health domain, users preferred the human metaphor (doctor) in terms of likability and enjoyability, while in the finance domain, no notable differences were observed between the human (financial advisor) and non-human (calculator) metaphors. Additionally, the choice of metaphor had no significant impact on users’ intentions to adopt the VUIs.
The use of fictional metaphors to inspire VUI design has long been of interest to HCI researchers. One of the most notable examples is drawing from the ”Computer” in Star Trek to inform VUI design principles (Turk, 2016). In a relevant study, Axtell & Munteanu (Axtell and Munteanu, 2021) conducted a conversation analysis comparing the dialogues of the ”Computer” in Star Trek: The Next Generation with interactions on Amazon Alexa. Their findings shed light on how VUIs could be designed to be more functional. Additionally, fictional metaphors can inspire a negative design heuristic, illustrating how not to design VUIs by highlighting ineffective or undesirable interaction patterns. This includes avoiding the dystopian, pervasive feeling of a ”Big Brother” presence from 1984 or the manipulative behavior exhibited by the Bene Gesserit Sisterhood through ”The Voice” in Dune (Feldman, 2024). Fictional metaphors can also inspire a sense of enchantment in VUI design, similar to how AI technologies often evoke a magical experience. This desired ”AI magic,” where users perceive advanced technology as almost supernatural, can be paralleled in VUIs by employing fictional metaphors that create a similar sense of wonder and fluid interaction. For instance, magic metaphors like using a ”magic wand” in GUIs to represent the undo command illustrate how these enchanting elements resonate with users (Lupetti and Murray-Rust, 2024), much like how wake-words in VUIs are sometimes compared to casting a spell (Desai and Twidale, 2023).
@@
Moreover, the issue of misaligned mental models, often described as the gulf between VUI capabilities and user expectations (Luger and Sellen, 2016), remains a key topic of ongoing research in the HCI community. One of the most pertinent approaches to studying this gulf is through the ”Mutual Theory of Mind” (MToM) Framework (Wang and Goel, 2022). This framework emphasizes that constructing a CUI’s understanding of how the user perceives the system—through analyzing communication cues—can be crucial in bridging this gap and aligning CUIs more effectively with users’ mental models. These communication cues include not only verbal signals such as language use and tone (Wang et al., 2021) but also non-verbal behaviors, like gestures, facial expressions, or pauses during interaction (Chan et al., 2024). In this paper, we take this idea further and propose designing conversational interfaces that adapt their presentational metaphor based on conversational contexts to ensure that the interface is aligned with the users’ mental models.
3. Toward Metaphor-Fluid Design
@@
The discussion so far has established that metaphors play an important role in VUIs. As discussed in the previous sections, metaphors help users construct mental models for interaction, guiding their expectations of system behavior (Carroll and Thomas, 1982; Carroll et al., 1988; Colburn and Shute, 2008). However, unlike in GUIs—where spatial and object-based metaphors provide a persistent conceptual framework (Blackwell, 2006)—metaphors in VUIs lack a stable visual anchor and must instead be inferred from interaction patterns. This fundamental difference has led to a design paradigm where system personas serve as implicit metaphors, mapping VUI behavior onto familiar social roles such as assistants, coaches, or companions (Desai and Chin, 2023; Jung et al., 2019; Desai et al., 2023a; Wang et al., 2020; Desai et al., 2023b; Lee et al., 2021; Motalebi et al., 2019). While these persona-based framings are intended to provide consistency, empirical findings suggest that users do not experience VUIs as singular, fixed entities. Instead, they reframe their mental models dynamically, adjusting their conceptualization of the system based on the nature of the interaction (Pradhan et al., 2019; Desai and Twidale, 2023). For example, Desai and Twidale (Desai and Twidale, 2023) found that when retrieving information, users tend to conceptualize VUIs as search engines or librarians, whereas in casual conversation, they are more likely to ascribe relational qualities, perceiving the VUI as a companion or conversational partner—sometimes within the same interaction session. These shifts occur fluidly and without explicit user effort, reflecting how people naturally adjust their conversational framing based on social and functional context.
@@
Despite this, current VUI design does not reflect this inherent metaphorical flexibility. Instead, designers have historically favored a single, unified persona, treating it as a stable interface identity that remains consistent across all interactions (Sadek et al., 2023). This approach assumes that users will develop trust and familiarity with a VUI if its persona is coherent and unchanging. However, prior research suggests that this assumption is not only misaligned with actual user expectations (Doyle et al., 2021; Desai et al., 2024a; Braun et al., 2019) but also overly simplistic (Desai et al., 2023a). Unlike GUIs, which can afford to maintain a static metaphorical structure, VUIs operate within a communicative medium where adaptability is already fundamental to human interaction. Individuals naturally adjust their language, tone, and conversational framing based on their conversational partner, the context of the exchange, and their intended outcomes. Imposing a single, inflexible metaphor onto a VUI is not only unnecessary but may actively conflict with how users intuitively engage with conversational systems. Furthermore, existing persona implementations rely on an inherent humanness metaphor, which introduces ethical concerns. The ”assistant” persona, for example, has been criticized for reinforcing gender stereotypes, highlighting the broader risks of anthropomorphic framing in conversational systems (Brahnam et al., 2011; Turk, 2016; Kuzminykh et al., 2020).
@@
In this paper, we introduce the concept of Metaphor-Fluid Design using VUIs as a central example. The primary aim is to present this novel design approach and examine its impact on user perception, particularly in comparison to traditional VUI implementations. To develop this approach, Study 1, guided by RQ1, investigates which metaphors naturally align with different use-contexts—commands, information-seeking, sociality, and error recovery. By identifying these patterns, Study 1 provides a foundation for designing VUIs that adapt their metaphorical framing to better align with user expectations in each context and explores the possibility of integrating non-human and fictional metaphors in VUI design. Study 2, guided by RQ2, then examines the implications of this approach by comparing a Metaphor-Fluid VUI to a Default VUI, assessing its effects on user perception, including enjoyment, likability, trust, intelligence, and intention to adopt. While prior work has shown that users already engage with and perceive VUIs in a metaphor-fluid manner, this study moves beyond observation to explore how metaphorical adaptation can be deliberately integrated into conversational systems and empirically evaluated. By expanding the metaphorical landscape beyond static, human-centric roles, we aim to demonstrate how dynamically adjusting metaphors in response to conversational context influences user perceptions and interaction experiences.

### v2

@@ Abstract.
Metaphors play a critical role in shaping user experiences with Voice User Interfaces (VUIs), yet existing designs rely on static, human-centric metaphors that fail to adapt to diverse contexts and user needs. We introduce Metaphor-Fluid Design, a novel approach that adjusts metaphorical representations based on conversational use-contexts. We compare this to a Default VUI, which models commercial VUIs commonly designed around the assistant persona, offering uniform interaction across contexts. In Study 1 (N=130), metaphors were mapped to four key use-contexts—commands, information seeking, sociality, and error recovery—along dimensions of formality and hierarchy, revealing distinct preferences for task-specific metaphorical designs. Study 2 (N=91) evaluates a Metaphor-Fluid VUI against a Default VUI, showing that Metaphor-Fluid Design enhances perceived intention to adopt, enjoyment, and likability. However, individual differences in metaphor preferences highlight the need for personalization. These findings challenge the one-size-fits-all paradigm and demonstrate Metaphor-Fluid Design’s potential for more adaptive human-AI interactions.
@@ 1. Introduction
Voice User Interfaces (VUIs) are becoming an integral part of daily interactions, embedded in smart speakers, mobile devices, automobiles, and wearable technologies. As these technologies proliferate, the ways in which users engage with and conceptualize them evolve dynamically. However, contemporary VUI design remains largely constrained by a single dominant metaphor—humanness (Doyle et al., 2019). Most commercial VUIs are designed to mimic human conversational abilities, using speech, turn-taking, and natural language processing to create an interaction model that feels intuitive (Edlund, 2019; Pradhan and Lazar, 2021). Within this overarching metaphor, VUIs are often framed as social entities that take on various roles, including teachers, therapists, coaches, advisors, and companions (Desai and Chin, 2023; Jung et al., 2019; Desai et al., 2023a; Wang et al., 2020; Desai et al., 2023b; Lee et al., 2021; Motalebi et al., 2019; Trajkova and Martin-Hammond, 2020; Purington et al., 2017). Among these, the assistant persona has emerged as the most prevalent, dominating commercial implementations and positioning the VUI as a helpful, subservient entity designed to execute user commands (Desai and Twidale, 2023; Doyle et al., 2021). While this metaphor provides a familiar interaction paradigm, it also imposes significant limitations. It assumes a fixed, one-size-fits-all identity for the VUI, failing to account for the fluid or context-dependent nature of human-VUI interactions across different tasks and contexts (Pradhan and Lazar, 2021; Desai et al., 2024b; Desai and Chin, 2023; Desai and Twidale, 2022). This rigidity leads to usability challenges (Cowan et al., 2017), mismatches between user expectations and system behavior (Doyle et al., 2021), and broader concerns related to identity surrounding anthropomorphic design choices (McMillan and Jaber, 2021; Desai and Twidale, 2022; Simpson and Crone, 2022; Pradhan and Lazar, 2021; Luger and Sellen, 2016; Turk, 2016).
@@
These challenges reflect a holistic issue in interface design: the role of metaphors in shaping user interaction. Metaphors have played a foundational role in shaping Human-Computer Interaction (HCI) (Neale and Carroll, 1997; Laurel, 1997; Laurel and Mountford, 1990; Carroll and Thomas, 1982; Carroll et al., 1988; Don et al., 1992). In Graphical User Interfaces (GUIs), they are structured into primary and secondary layers: the desktop metaphor, for instance, serves as the primary organizing concept, with secondary metaphors—such as files, folders, and recycling bins—providing finer interaction affordances (Blackwell, 2006). These layered metaphors help users form coherent mental models of system functionality. In contrast, VUIs lack a stable visual representation, making their metaphorical framing less explicit. Instead, this framing is operationalized through persona design, where a VUI’s voice, linguistic style, and response behaviors collectively establish its identity (Pradhan and Lazar, 2021). Unlike GUIs, where metaphors map interactions onto familiar objects, VUIs inherently imply humanness, reinforcing anthropomorphic interpretations. The presence of human-like speech cues, names, and conversational norms encourages users to assign social roles (e.g., therapist or companion) to VUIs, often expecting them to behave with intelligence, adaptability, and even emotional awareness (Luger and Sellen, 2016; Sutton et al., 2019).
@@
Despite this anthropomorphic framing, users do not engage with VUIs as static entities (Pradhan et al., 2019). Instead, they shift their metaphorical framings depending on use-contexts—the functional settings in which interactions take place (Desai and Twidale, 2023). Prior research has identified several recurring use-contexts in VUI interactions, including commands, where users expect efficiency and compliance (e.g., ”Turn off the lights”); information seeking, where they seek structured knowledge retrieval (e.g., ”What is the capital of Japan?”); sociality, where they engage in casual conversation (e.g., ”Tell me a joke”); and error recovery, where they attempt to correct or refine system responses (e.g., ”That’s not what I meant—try again”) (Sciuto et al., 2018; Desai and Chin, 2023; Ammari et al., 2019; Bentley et al., 2018). In each of these use-contexts, users adopt distinct metaphorical descriptions: they conceptualize a VUI as a butler or secretary when issuing commands, as an expert librarian when retrieving information, and as a companion in social exchanges (Desai and Twidale, 2023). However, commercial VUIs impose a singular, unchanging persona across all contexts, failing to adapt to the shifting expectations users bring to different types of interactions. This rigidity results in fundamental misalignments—when users anticipate authoritative expertise but receive generic, non-committal responses, or when they expect a social exchange but encounter the same neutral, task-oriented demeanor, the interaction feels unnatural and disengaging (Desai et al., 2023a). These breakdowns not only present usability challenges but also raise deeper ethical concerns.
@@
Building on prior work, we introduce and propose a shift toward Metaphor-Fluid Design (as illustrated in Figure 1), a novel design approach that moves beyond static persona models by enabling VUIs to dynamically shift their metaphorical framing based on conversational context. Rather than adhering to a singular metaphor—such as the assistant—this approach allows the system to adopt multiple personas, aligning its interaction style with the expectations of the task at hand. Crucially, Metaphor-Fluid Design also incorporates non-human and fictional metaphors, which have been shown to be both relevant and frequently used by users (Desai et al., 2024b; Desai and Twidale, 2023; Jung et al., 2022). While commercial VUIs default to anthropomorphic personas, users often describe them in non-human terms, likening them to machines, calculators, encyclopedias, or even mythical entities like genies or ghosts (Desai and Twidale, 2023). By broadening the metaphorical landscape beyond human roles, this approach creates greater flexibility in interaction design, allowing VUIs to respond more intuitively to diverse user expectations. This work is guided by two primary research questions:
@@
These questions are explored through two studies. The first study (N=130) maps user-generated metaphors to different conversational contexts, revealing systematic variations in metaphorical preferences across tasks. Findings show that users do not conceptualize VUIs through a singular metaphor but instead shift their framings based on the use-context. The second study (N=91) empirically evaluates a Metaphor-Fluid VUI—which adapts its metaphorical framing dynamically—against a Default VUI that consistently operationalizes the assistant metaphor. Results indicate that Metaphor-Fluid design leads to significantly higher perceived likability, enjoyment, and intention to adopt, showing that aligning a VUI’s persona with context-specific expectations positively impacts user perceptions. However, findings also highlight individual differences in metaphor preferences.
@@ 2.3. Metaphors in VUIs
    Simplistic Representations: While users develop complex and context-dependent mental models of VUIs that change with the type of interaction or domain of use, designers often rely on a single overarching metaphor to represent these systems. This singular metaphorical framing may fail to capture the nuanced ways in which users interact with VUIs across different contexts (Desai and Twidale, 2022). For example, in a user study involving older adults interacting with a VUI designed as an ”exercise coach,” participants expressed a preference for the VUI to also behave like a ”friend” when initiating the session, checking on their progress, and providing motivation (Desai et al., 2023a). The complexity of user expectations and the diverse roles VUIs play are not adequately addressed by a one-size-fits-all metaphor, which can lead to a mismatch between what users expect and what the system delivers, ultimately impacting usability and user satisfaction. Moreover, users’ perceptions of VUIs are also highly fluid, shifting depending on the context and specific interactions. Pradhan et al. (Pradhan and Lazar, 2021) found that users alternately view VUIs as either a ”person” or a ”thing,” or something in between, depending on the specific task or setting.
@@
    Misaligned Mental Models: Misaligned mental models arise when the human-like metaphors underlying VUIs heighten user expectations, leading to either over or under-calibration of trust. Users may overestimate the VUI’s abilities, assuming it possesses human-like cognitive and conversational skills, or underestimate its utility due to skepticism about its capabilities (Cowan et al., 2017). This misalignment of expectations impacts trust calibration—users may become overly reliant on the VUI, expecting it to perform beyond its technical limitations, or conversely, they may dismiss its utility prematurely (Luger and Sellen, 2016). These fluctuating expectations can result in inconsistent or infrequent use (Cowan et al., 2017) and, eventually, complete abandonment of the technology (Trajkova and Martin-Hammond, 2020).
The evolving landscape of VUI design reflects a broader understanding of the complexities involved in metaphor selection.
2.3.1. Mitigating Challenges in VUI Design
@@
As the VUI design space evolves, various approaches have been proposed to address persona design challenges. One promising direction involves developing VUIs with non-human roles. Desai and Twidale (Desai and Twidale, 2023) developed a framework for metaphor contextualization in VUIs, categorizing metaphor use along four dimensions: type (human, non-human, fictional), mentioned by (users, designers, manufacturers, researchers), knowledge drawn upon, and motivation for use. While human metaphors dominate VUI design and research, non-human and fictional metaphors were equally popular among users for sense-making, highlighting a dichotomy between designer presentations and user interpretations that points to the need for a ”pragmatic approach” (Carroll et al., 1988) to explore the design space between system capabilities and user expectations.
@@
Recent CUI literature has explored non-human conversational agents. Jung et al. (Jung et al., 2022) used the Great Chain of Being framework [^2] A theoretical framework that organizes the world into a hierarchical structure using metaphors of God, humans, animals, plants, and inanimate objects. to investigate how non-human metaphors (God, Animal, Plant, Book) compared to human metaphors in chatbot perceptions, finding no significant preference for human metaphors. Similarly, VUI studies examining human versus non-human metaphors found domain-specific preferences: users favored human metaphors in health contexts but showed no clear preference in finance (Desai et al., 2024a).
Fictional metaphors have long interested HCI researchers, notably through Star Trek’s ”Computer” (Turk, 2016). Axtell and Munteanu (Axtell and Munteanu, 2021) compared Star Trek dialogues with Amazon Alexa interactions to explore functional design principles. Fictional metaphors can also provide negative design heuristics, highlighting undesirable patterns like the dystopian ”Big Brother” from 1984 or manipulative ”Voice” from Dune (Feldman, 2024), while inspiring enchantment through ”AI magic” experiences (Lupetti and Murray-Rust, 2024; Desai and Twidale, 2023).
@@
The persistent issue of misaligned mental models—the gulf between VUI capabilities and user expectations (Luger and Sellen, 2016)—remains central to HCI research. Two pertinent approaches address this through the Mutual Theory of Mind framework (Wang and Goel, 2022) and partner models (Doyle et al., 2021). The MToM framework emphasizes analyzing communication cues (Wang et al., 2021; Chan et al., 2024) to align CUIs with user mental models, while partner models represent users’ cognitive beliefs about VUI communicative capabilities—multi-dimensional perceptions of competence, dependability, human likeness, and cognitive flexibility that update dynamically through interaction (Doyle et al., 2021; Gessinger et al., 2025). Building on these concepts, we propose designing conversational interfaces that adapt their presentational metaphor based on conversational contexts to ensure alignment with users’ mental models.
@@
Metaphors play an important role in VUIs by helping users construct mental models for interaction and guiding expectations of system behavior (Carroll and Thomas, 1982; Carroll et al., 1988; Colburn and Shute, 2008). As discussed in the previous sections, unlike GUIs where spatial and object-based metaphors provide persistent visual frameworks (Blackwell, 2006), VUI metaphors lack stable visual anchors and must be inferred from interaction patterns. This fundamental difference has led to a design paradigm where system personas serve as implicit metaphors, mapping VUI behavior onto familiar social roles such as assistants, coaches, or companions (Desai and Chin, 2023; Jung et al., 2019; Desai et al., 2023a; Wang et al., 2020; Desai et al., 2023b; Lee et al., 2021; Motalebi et al., 2019; Desai and Chin, 2021). Empirical findings reveal that users do not experience VUIs as singular, fixed entities but instead dynamically reframe their mental models based on interaction context (Pradhan et al., 2019; Desai and Twidale, 2023). In a previous study, researchers found that users conceptualize VUIs as search engines or librarians when retrieving information, but as companions during casual conversation—sometimes within the same session (Desai and Twidale, 2023). These shifts occur fluidly, reflecting how people naturally adjust conversational framing based on social and functional context.
@@
Current VUI design does not reflect this inherent flexibility as conversation designers have historically favored single, unified personas that remain consistent across all interactions (Sadek et al., 2023), assuming users will develop trust through coherent, unchanging identities. However, this assumption is misaligned with user expectations (Doyle et al., 2021; Desai et al., 2024b; Braun et al., 2019) and overly simplistic (Desai et al., 2023a). VUIs operate within a communicative medium where adaptability is fundamental to human interaction. Imposing a single, inflexible metaphor may actively conflict with how users intuitively engage with conversational systems and reinforces problematic anthropomorphic framing, as seen in the ”assistant” persona’s reinforcement of gender stereotypes (Brahnam et al., 2011; Turk, 2016; Kuzminykh et al., 2020).
@@
In this paper, we introduce the concept of Metaphor-Fluid Design using VUIs as a central example. The primary aim is to present this novel design approach and examine its impact on user perception, particularly in comparison to traditional VUI implementations.